# Xiaofeng Yang, Yuhui Zhang, Yifeng Zheng, Xiuqi Huang, Minfeng Zhuparallel-$\parallel$, Yuxin Ma, Bin Cui§, Wei Chen

### v2

@@ VIII Related Work
LLM-based Data Analysis. LLMs have demonstrated significant capabilities in semantic understanding and logical reasoning, facilitating complex data analysis tasks within conversational interfaces [6, 22]. Early empirical studies [76] investigate how data analysts interact with LLMs and identify challenges like contextual data retrieval, prompt refinement, and code adaptation. Table-GPT [77] fine-tunes LLMs on synthesized table-task data to enhance their table-understanding abilities. Inspired by Chain-of-Thought prompting [43], Chat2Query [23] decomposes the NL2SQL task into multiple steps to improve generation quality. Similarly, InsightPilot [78] automates the discovery of data insights and synthesizes them into high-level overviews. Chat2Data [79] leverages domain knowledge based on vector databases to mitigate LLMs’ hallucination issues. While most existing works focus on individual tasks and generate results in an end-to-end manner, we introduce a unified platform for various data tasks in a computational notebook workspace. It supports user intervention of intermediate results, enabling more flexible human-LLM collaboration for data analysis in BI scenarios.
IX Conclusion
@@
In this paper, we present DataLab, a unified platform for business intelligence that combines an LLM-based agent framework with a computational notebook interface. DataLab features a domain knowledge incorporation module, an inter-agent communication mechanism, and a cell-based context management strategy. These components enable seamless integration of LLM assistance with user customization, making DataLab well-suited for practical BI scenarios. Extensive experiments on both research benchmarks and real-world business datasets demonstrate the effectiveness of DataLab.

### v3

@@ IX Related Work
LLM-based Data Analysis. LLMs have shown remarkable abilities in semantic understanding and logical reasoning, enabling complex data analysis through conversational interfaces [6, 22, 69]. For example, Table-GPT [70] fine-tunes LLMs on synthesized table-task data to enhance their table-understanding abilities. Chat2Query [23] decomposes NL2SQL tasks into multiple steps to improve generation quality. InsightPilot [71] automates the discovery of data insights and synthesizes them into high-level overviews. Moreover, Chat2Data [15] leverages domain knowledge based on vector databases to mitigate LLMs’ hallucination issues, while our work uniquely facilitates domain-specific data analysis by automatically extracting knowledge from enterprise scripts and data lineage information, avoiding manual curation. Unlike existing works that focus on end-to-end results, we introduce a novel notebook interface that supports flexible human intervention - allowing refinement of intermediate SQL, Python, or charts - while unifying tasks fragmented across prior task-specific approaches like LIDA [30] and PURPLE [21].
X Conclusion
@@
This paper introduces DataLab, a unified BI platform that combines an LLM-based agent framework with a notebook interface. DataLab features a domain knowledge incorporation module, an inter-agent communication mechanism, and a cell-based context management strategy. These components enable seamless integration of LLM assistance with user customization, making DataLab well-suited for practical BI scenarios. DataLab has proven effective on both research benchmarks and real-world business datasets from Tencent.